OpenStudy (anonymous):

One number is 2 times as large as another. The sum of their reciprocals is 15/2. Find the two numbers. (smaller value) (larger value)

OpenStudy (anonymous):

Let smaller number be S Let larger number be L L = 2*S 1/L + 1/S = 15/2 Solve the equations.

OpenStudy (anonymous):

1/2S + 1/S = 15/2 3/2S = 15/2 2S = 3*2/15 S = 3/15 = 1/5 L = 2/5